This spinner is spun 36 times. The spinner landed on A 6 times, on B 21 times, and on C 9 times. Compute the empirical probability that the spinner will land on B.
Please help me!!Thanks a lot

Answer by Theo(13342)

spinner is spun 36 times.
it lands on A 6 times
it lands on B 21 times
it lands on C 9 times.

the empirical probability that it will land on A is equal to 6/36.
the empirical probability that it will land on B is equal to 21/36.
the empirical probability that it will land on C is equal to 9/36.

your answer is:

the empirical probability that it will land on B is equal to 21/36.

empirical means "from experience".

if something happens x% of the time from past experience, then you can say that the probability that it will occur again is x%.

that's exactly what we did above.

the x was 21
the total was 36.

x% would be 21/36 * 100 = 58.3% of the time.

the total percent should be equal to 100%.

you have:

p(A) = 6/36 = 16.7%
p(B) = 21/36 = 58.3%
p(C) = 9/36 = 25%

add them all up together and you get 16.7 + 58.3 + 25 = 100%